OpenStudy (anonymous):

Three brothers spent different amounts of money on their recent vacation. Matt spent $25 less than Todd. Jeff spent 3 times as much as Matt. How much did each brother spend if they spent a total of $275?

OpenStudy (anonymous):

T = M + T + J = M + ( M + 5) + 3M = 275 @magicbananas Can you continue?

OpenStudy (anonymous):

how to find perimeter of a shape?

OpenStudy (anonymous):

can you be more elaborate?

OpenStudy (anonymous):

i know but i dont understand this at all....

OpenStudy (anonymous):

"Jeff spent 3 times as much as Matt" J = 3 * M Do you understand it?

OpenStudy (anonymous):

kind of...

OpenStudy (anonymous):

not really nevermind........